Mechanics

A car is moving in a circular path of radius 100 m with velocity of 200 m/sec such that in each sec its velocity increases by 100 m/s, the net acceleration of car will be - (in m/sec)

Anirban Jyoti
7 Years agoGrade

1 Answer

Arun
7 Years ago
 
There are 2 components of acceleration.
  1. Towards center: V2/R = 200 * 200/100 = 20 m/s2
  2. Tangent( orthogonal to center) = 100 m/s2
Resultant of above = sqrt (10000 + 400) = sqrt (10400)